# Singular dynamics for discrete weak K.A.M. solutions of exact twist maps

### v1

@@ 3.1. $\Sigma_{+}$ is well-defined, non-decreasing and continuous
If $u\in C^{0}(\mathbb{T})$, then $\Sigma_{+}$ is non-decreasing. That is, if $x_{1}<x_{2}$, then $\Sigma_{+}(x_{1})\leq\Sigma_{+}(x_{2})$.
@@
Suppose that $x_{1}<x_{2}$ but $y_{1}>y_{2}$, where $y_{1}\in\Sigma_{+}(x_{1})$ and $y_{2}\in\Sigma_{+}(x_{2})$. We will use Non-crossing Lemma (Lemma 2.19) to reduce a contradiction. We know $(x_{1}-x_{2})(y_{1}-y_{2})<0$ and the Non-crossing Lemma says
$S(x_{1},y_{1})+S(x_{2},y_{2})>S(x_{1},y_{2})+S(x_{2},y_{1}).$
Using the definition of the set-valued map $\Sigma_{+}$, we have that
$u(y_{1})-S(x_{1},y_{1})=T^{+}[u](x_{1})=\sup_{y\in\mathbb{R}}\big{\{}u(y)-S(x_{1},y)\big{\}}\geq u(y_{2})-S(x_{1},y_{2});$
$u(y_{2})-S(x_{2},y_{2})=T^{+}[u](x_{2})=\sup_{y\in\mathbb{R}}\big{\{}u(y)-S(x_{2},y)\big{\}}\geq u(y_{1})-S(x_{2},y_{1}).$
@@
This contradicts the Non-crossing Lemma and we complete the proof. ∎
@@
If $u$ is a discrete weak K.A.M. solution, then
$\bigcup_{x\in\mathbb{R}}\Sigma_{+}(x)=\mathbb{R}.$
Indeed, for any $y\in\mathbb{R}$, there exists $x\in\mathbb{R}$ such that $(x,y)\in\mathcal{NS}(u)$. Specifically, $y\in\Sigma_{+}(x)$.
@@
Thanks to Remark 2.4, for any $y\in\mathbb{R}$, there exists $x\in\mathbb{R}$ such that
$T_{-}[u](y)=u(x)+S(x,y).$
Since $u$ is a discrete weak K.A.M. solution, we have
$u(y)+\bar{S}=T_{-}[u](y)=u(x)+S(x,y).$
That is, the pair $(x,y)$ is $u$-calibrated. Hence by Remark 2.11, we have $T^{+}[u](x)=u(y)-S(x,y)$ That is, $y\in\Sigma_{+}(x)$. ∎
@@
If $u$ is a discrete weak K.A.M. solution, then for any $x\in\mathbb{R}$, the set $\Sigma_{+}(x)$ is a singleton.
@@
Suppose that there exists $y_{1}<y_{2}$ such that $y_{1},y_{2}\in\Sigma_{+}(x_{0})$. By Lemma 3.1, for any $x>x_{0}$, we have $\Sigma_{+}(x)\geq y_{2}$; for any $x<x_{0}$, we have $\Sigma_{+}(x)\leq y_{1}$. Hence by Lemma 3.2, we have
$(y_{1},y_{2})\subset\Sigma_{+}(x_{0}).$
That is,
$u(y)-S(x_{0},y)=T^{+}[u](x_{0}),\quad\forall y\in(y_{1},y_{2}).$
Since $S$ is $C^{1}$, we have that $u$ is differentiable on $(y_{1},y_{2})$ and
$\ \mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y)=\partial_{2}S(x_{0},y),\quad\forall y\in(y_{1},y_{2}).$
Using Theorem 2.15, we have
$(x_{0},-\partial_{1}S(x_{0},y))=F^{-1}(y,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}y)\in\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u),\quad\forall y\in(y_{1},y_{2}).$
Hence
$-\partial_{1}S(x_{0},y)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{0}),\quad\forall y\in(y_{1},y_{2}),$
which contradicts $y_{1}\neq y_{2}$ and that $S$ is ferromagnetic. ∎
Using (b) of Lemma 2.6, we have the following corollary.
Corollary 2.
If $u$ is a discrete weak K.A.M. solution, then $T^{+}[u]$ is differentiable everywhere.
If $u$ is a discrete weak K.A.M. solution, then $\Sigma_{+}:\mathbb{R}\rightarrow\mathbb{R}$ is a well-defined map. It follows from Lemma 3.1 that $\Sigma_{+}$ is non-decreasing as a real-valued map. Since discontinuities of monotonic function are jump discontinuities, but Lemma 3.2 implies that $\Sigma_{+}$ has no jump discontinuity. We conclude that $\Sigma_{+}$ is continuous.
Utilizing the periodicity of $u$ and $T^{+}[u]$, along with the diagonal-periodicity of $S$, we have
$y=\Sigma_{+}(x)$ $\quad\Longleftrightarrow\quad T^{+}[u](x)=u(y)-S(x,y)$
$\quad\Longleftrightarrow\quad T^{+}[u](x+1)=u(y+1)-S(x+1,y+1)$
$\quad\Longleftrightarrow\quad y+1=\Sigma_{+}(x+1).$
This implies $\Sigma_{+}(x+1)=\Sigma_{+}(x)+1$. In other words, $\Sigma_{+}$ is a lift of a circle map of degree $1$.
Remark 3.4.
Notice that Zavidovique’s notes [Zav23, Corollary 5.3.13] states the regularity of $T^{+}[u]$ whenever $u$ is a weak K.A.M. solution.
Proposition 3.5 ([Zav23]).
If $u$ is a discrete weak K.A.M. solution, then $T^{+}[u]$ is $C^{1}$.
Our method is different from Zavidovique’s because Zavidovique used one theorem due to Arnaud (see [Arn11, Corollary 2]). Furthermore, we will strengthen Zavidovique’s result in Corollary 4 to $T^{+}[u]$ is $C^{1,1}$, provided $u$ is a discrete weak K.A.M. solution.
3.2. $\Sigma_{+}$ is Lipschitz
First we prove a strengthened version of Lemma 3.2.
@@ 3.2. $\Sigma_{+}$ is Lipschitz
Let $u$ be a discrete weak K.A.M. solution. For any $y\in\mathbb{R}$, the preimage $(\Sigma_{+})^{-1}(y)$ is either a singleton or a non-degenerate closed bounded interval. If $(\Sigma_{+})^{-1}(y)$ is a singleton, denoted by $x$, then $(x,y)\in\mathcal{NS}(u)$. If $(\Sigma_{+})^{-1}(y)$ is a non-degenerate closed bounded interval, denoted by $[x_{-},x_{+}]$, then $(x_{-},y),(x_{+},y)\in\mathcal{NS}(u)$.
@@
By Lemma 3.2, we know that $(\Sigma_{+})^{-1}(y)$ is nonempty for all $y\in\mathbb{R}$.
Case 1: $(\Sigma_{+})^{-1}(y)$ is a singleton $x$. By Lemma 3.2, we can find $x^{\prime}\in\mathbb{R}$ such that $(x^{\prime},y)\in\mathcal{NS}(u)$ and $x^{\prime}\in(\Sigma_{+})^{-1}(y)$. Hence by the singleton property of $(\Sigma_{+})^{-1}(y)$, we have $x^{\prime}=x$ and $(x,y)\in\mathcal{NS}(u)$.
Case 2: $(\Sigma_{+})^{-1}(y)$ contains at least two different points $x_{1}<x_{2}$. Since $\Sigma_{+}$ is continuous and satisfies $\Sigma_{+}(x+1)=\Sigma_{+}(x)+1$, we claim that
$x_{-}:=\inf\{x\in\mathbb{R}\mid\Sigma_{+}(x)=y\},\quad x_{+}:=\sup\{x\in\mathbb{R}\mid\Sigma_{+}(x)=y\}$
satisfy $-\infty<x_{-}\leq x_{1}<x_{2}\leq x_{+}<+\infty$ and $x_{-},x_{+}\in(\Sigma_{+})^{-1}(y)$. Since $\Sigma_{+}$ is non-decreasing, we have $(\Sigma_{+})^{-1}(y)=[x_{-},x_{+}]$, which is a non-degenerate closed bounded interval.
Before showing $(x_{-},y),(x_{+},y)\in\mathcal{NS}(u)$, let us first prove that the cardinal of
[…]
is at most countable. We give an injection from $\mathcal{C}$ to $\mathbb{Q}$. We consider the map
$(\Sigma_{+})^{-1}:\mathcal{C}$ $\rightarrow 2^{\mathbb{R}}$
$y$ $\mapsto[x_{-}(y),x_{+}(y)]$
Since $\Sigma_{+}$ is a well-defined map, for $y_{1}\neq y_{2}$, we have
$[x_{-}(y_{1}),x_{+}(y_{1})]\cap[x_{-}(y_{2}),x_{+}(y_{2})]=\emptyset.$
By Axiom of Choice, we can pick a rational number $q(y)$ from each $[x_{-}(y),x_{+}(y)]$. Hence for $y_{1}\neq y_{2}$, we get $q(y_{1})\neq q(y_{2})$. This defines an injection from $\mathcal{C}$ to $\mathbb{Q}$, whose cardinal is countable.
If $(\Sigma_{+})^{-1}(y)$ is a non-degenerate closed bounded interval $[x_{-},x_{+}]$, then we can find an increasing sequence $\{x_{n}\}_{n\in\mathbb{Z}_{+}}$ such that $x_{n}\rightarrow x_{-}$ and
$(\Sigma_{+})^{-1}(y_{n}),\quad\text{ where }y_{n}=\Sigma_{+}(x_{n}),$
is a singleton for all $n\in\mathbb{Z}_{+}$. Suppose this is incorrect, then there exists $\delta>0$ such that for all $x\in[x_{-}-\delta,x_{-}]$, $(\Sigma_{+})^{-1}(\Sigma_{+}(x))$ is not a singleton (so it must be a non-degenerate closed interval). Since $x_{-}=\inf\{x\mid\Sigma_{+}(x)=y\}$, we have $\Sigma_{+}(x_{-}-\delta)<y$. Now using the fact that $\mathcal{C}$ is at most countable, we claim there exists $y^{\prime}$ such that $\Sigma_{+}(x_{-}-\delta)<y^{\prime}<y$ and $(\Sigma_{+})^{-1}(y^{\prime})$ is a singleton, which is denoted by $x^{\prime}$. Since $\Sigma_{+}$ is non-decreasing, we have $x^{\prime}\in[x_{-}-\delta,x_{-}]$, but $(\Sigma_{+})^{-1}(y^{\prime})$ is a singleton. Contradiction!
Thus by Case 1, the sequence $\{x_{n}\}_{n\in\mathbb{Z}_{+}}$ satisfies $(x_{n},y_{n})\in\mathcal{NS}(u)$. Since $\mathcal{NS}(u)$ is closed and $\lim_{n\to+\infty}(x_{n},y_{n})=(x_{-},y)$, we have $(x_{-},y)\in\mathcal{NS}(u)$. By a similar argument, we similarly obtain $(x_{+},y)\in\mathcal{NS}(u)$. ∎
Corollary 3.
Let $u$ be a discrete weak K.A.M. solution. For any $x\in\mathbb{R}$, there exist $x^{\prime},x^{\prime\prime}\in\mathbb{R}$ such that $x^{\prime}\leq x\leq x^{\prime\prime}$, $\Sigma_{+}(x^{\prime})=\Sigma_{+}(x)=\Sigma_{+}(x^{\prime\prime})$ and $x^{\prime},x^{\prime\prime}\in\pi_{1}(\mathcal{NS}(u))$.

### v2

@@ 3.1. $\Sigma_{+}[u]$ is a non-decreasing continuous real-valued mapping
If $u\in C^{0}(\mathbb{T})$, then $\Sigma_{+}[u]$ is non-decreasing. In other words, if $x_{1}<x_{2}$, then $\Sigma_{+}[u](x_{1})\leq\Sigma_{+}[u](x_{2})$.
@@
Suppose that $x_{1}<x_{2}$ but $y_{1}>y_{2}$, where $y_{1}\in\Sigma_{+}(x_{1})$ and $y_{2}\in\Sigma_{+}(x_{2})$. We will use the Non-crossing Lemma (see Hypothesis 3) to reduce a contradiction. We know $(x_{1}-x_{2})(y_{1}-y_{2})<0$ and the Non-crossing Lemma says
$S(x_{1},y_{2})+S(x_{2},y_{1})<S(x_{1},y_{1})+S(x_{2},y_{2}).$
Using the definition of the optimal forward mapping $\Sigma_{+}[u]$, from $y_{1}\in\Sigma_{+}(x_{1})$ and $y_{2}\in\Sigma_{+}(x_{2})$, we get that
$u(y_{1})-S(x_{1},y_{1})\geq u(y_{2})-S(x_{1},y_{2});$
$u(y_{2})-S(x_{2},y_{2})\geq u(y_{1})-S(x_{2},y_{1}).$
@@
This contradicts the Non-crossing Lemma and thus we complete the proof. ∎
@@
If $u=T^{-}[v]$, $v\in SC(\mathbb{T})$, then for any $x\in\mathbb{R}$, the set $\Sigma_{+}[u](x)$ is a singleton.
@@
Suppose that there exists $y_{1}<y_{2}$ such that $y_{1},y_{2}\in\Sigma_{+}[u](x_{0})$. By Lemma 3.1, for any $x>x_{0}$, we have $\Sigma_{+}[u](x)\geq y_{2}$ and for any $x<x_{0}$, we have $\Sigma_{+}[u](x)\leq y_{1}$. In other words, if $\Sigma_{+}[u](x)\cap(y_{1},y_{2})\neq\emptyset$, then $x=x_{0}$. Now, take $z_{1},z_{2}$ such that $y_{1}<z_{1}<z_{2}<y_{2}$. For any $x\in\Sigma_{-}[v](z_{i})$, $i=1,2$, by Lemma 2.5, we have $z_{i}\in\Sigma_{+}[u](x)\cap(y_{1},y_{2})$, which implies $x=x_{0}$. Thus, $x_{0}\in\Sigma_{-}[v](z_{1})\cap\Sigma_{-}[v](z_{2})$. By Lemma 2.20, since $v$ is semiconcave, we must have $z_{1}=z_{2}$. This leads to a contradiction. ∎
If $u=T^{-}[v]$, $v\in SC(\mathbb{T})$, then $\Sigma_{+}[u]:\mathbb{R}\rightarrow\mathbb{R}$ is a non-decreasing real-valued mapping. Since discontinuities of monotonic function are jump discontinuities, but Lemma 2.6 implies that $\Sigma_{+}[u]$ has no jump discontinuity, we conclude that $\Sigma_{+}[u]$ is continuous for any $u\in T^{-}(C^{0}(\mathbb{T}))$. Utilizing the periodicity of $u$, along with the diagonal-periodicity of $S$, we deduce from the definition that $\Sigma_{+}[u](x+1)=\Sigma_{+}[u](x)+1$. In other words, $\Sigma_{+}[u]$ is a lift of a circle map (not necessarily invertible) of degree $1$. We conclude that
@@
If $u=T^{-}[v]$, $v\in SC(\mathbb{T})$, then $\Sigma_{+}[u]$ is a non-decreasing continuous real-valued mapping, which is a lift of a circle map of degree $1$.∎
3.2. $\Sigma_{+}[u]$ is Lipschitz continuous
Let $u=T^{-}[v]$, where $v\in SC(\mathbb{T})$. It has already been established that $\Sigma_{+}[u]$ is Lipschitz continuous on the subset $\mathcal{O}[v]=\Sigma_{-}[v](\mathbb{R})\subset\mathbb{R}$ (see Lemma 2.23). To prove $\Sigma_{+}[u]$ is Lipschitz continuous on the entire $\mathbb{R}$, we proceed as follows:
For any $x\in\mathbb{R}$, we will identify another point $x^{\prime}\in\Sigma_{-}[v](\mathbb{R})$ such that $\Sigma_{+}[u](x)=\Sigma_{+}[u](x^{\prime})$. Then, by leveraging the established Lipschitz property of $\Sigma_{+}[u]$ on $\Sigma_{-}[v](\mathbb{R})$, we will deduce the Lipschitz continuity of $\Sigma_{+}[u]$ at $x$.
Lemma 3.4.
Let $u=T^{-}[v]$, $v\in SC(\mathbb{T})$. For any $y\in\mathbb{R}$, the preimage $(\Sigma_{+}[u])^{-1}(y)$ is either a singleton or a non-degenerate closed bounded interval. If $(\Sigma_{+}[u])^{-1}(y)$ is a singleton, denoted by $x$, then $x\in\Sigma_{-}[v](y)$. If $(\Sigma_{+}[u])^{-1}(y)$ is a non-degenerate closed bounded interval, denoted by $[x_{-},x_{+}]$, then $x_{-},x_{+}\in\Sigma_{-}[v](y)$.
@@ 3.2. $\Sigma_{+}[u]$ is Lipschitz continuous
By Lemma 2.6, we know that for any $y\in\mathbb{R}$, the set $(\Sigma_{+}[u])^{-1}(y)$ is nonempty.
Case 1: If $(\Sigma_{+}[u])^{-1}(y)$ is a singleton $\{x\}$. By Remark 2.4 (a), we can find $x^{\prime}\in\mathbb{R}$ such that $x^{\prime}\in\Sigma_{-}[u](y)$. Then by Lemma 2.5, $x^{\prime}\in(\Sigma_{+}[u])^{-1}(y)$. Since $(\Sigma_{+}[u])^{-1}(y)$ is a singleton, we must have $x^{\prime}=x$, implying $x\in\Sigma_{-}[u](y)$.
Case 2: If $(\Sigma_{+}[u])^{-1}(y)$ contains at least two distinct points $x_{1}<x_{2}$. Since $\Sigma_{+}[u]$ is continuous and satisfies $\Sigma_{+}[u](x+1)=\Sigma_{+}[u](x)+1$, we define
$x_{-}:=\inf\{x\in\mathbb{R}\mid\Sigma_{+}[u](x)=y\},\quad\quad x_{+}:=\sup\{x\in\mathbb{R}\mid\Sigma_{+}[u](x)=y\}.$
Clearly, $-\infty<x_{-}\leq x_{1}<x_{2}\leq x_{+}<+\infty$ and $x_{-},x_{+}\in(\Sigma_{+}[u])^{-1}(y)$. Since $\Sigma_{+}[u]$ is non-decreasing, $(\Sigma_{+}[u])^{-1}(y)=[x_{-},x_{+}]$, which is a non-degenerate closed bounded interval.
To prove $x_{-},x_{+}\in\Sigma_{-}[v](y)$, we first establish that the cardinality of
[…]
is at most countable. Define a mapping
$(\Sigma_{+}[u])^{-1}:\mathcal{C}$ $\rightarrow 2^{\mathbb{R}}$
$y$ $\mapsto[x_{-}(y),x_{+}(y)].$
For $y_{1}\neq y_{2}$, the intervals $[x_{-}(y_{1}),x_{+}(y_{1})]$ and $[x_{-}(y_{2}),x_{+}(y_{2})]$ are disjoint. Using the Axiom of Choice, we can select a rational number $q(y)$ from each $[x_{-}(y),x_{+}(y)]$, thus defining an injection from $\mathcal{C}$ into $\mathbb{Q}$.Since $\mathbb{Q}$ is countable, $\mathcal{C}$ is also at most countable.
Now, we show $x_{-},x_{+}\in\Sigma_{-}[v](y)$. Suppose $(\Sigma_{+}[u])^{-1}(y)=[x_{-},x_{+}]$ is a non-degenerate interval. We claim there exists a sequence $\{x_{n}\}_{n\in\mathbb{Z}_{+}}$ with $x_{n}\rightarrow x_{-}$ and $(\Sigma_{+}[u])^{-1}(\Sigma_{+}[u](x_{n}))$ is a singleton for each $n$. Suppose this is not true, then there exists $\delta>0$ such that $(\Sigma_{+}[u])^{-1}(\Sigma_{+}[u](x))$ is not a singleton for all $x\in[x_{-}-\delta,x_{-}]$. Since $x_{-}=\inf\{x\mid\Sigma_{+}[u](x)=y\}$, we have $\Sigma_{+}[u](x_{-}-\delta)<y$. Using the countability of $\mathcal{C}$, there exists $y^{\prime}$ such that $\Sigma_{+}(x_{-}-\delta)<y^{\prime}<y$ and $(\Sigma_{+}[u])^{-1}(y^{\prime})$ is a singleton, say $x^{\prime}$. Then since $\Sigma_{+}$ is non-decreasing, $x^{\prime}\in[x_{-}-\delta,x_{-}]$, which contradicts the assumption that $(\Sigma_{+}[u])^{-1}(y^{\prime})$ is not a singleton.
Define $y_{n}:=\Sigma_{+}[u](x_{n})$, and thus $x_{n}\in\Sigma_{-}[v](y_{n})$ for all $n$. Define $f_{n}(x):=v(x)+S(x,y_{n})$ and $f(x):=v(x)+S(x,y)$. Since $f_{n}\rightarrow f$ uniformly on compact sets, $\inf f_{n}$ converges to $\inf f$. Note that $\inf f_{n}=v(x_{n})+S(x_{n},y_{n})$, so
$\inf f=\lim_{n}\inf f_{n}=\lim_{n}\left(v(x_{n})+S(x_{n},y_{n})\right)=v(x_{-})+S(x_{-},y).$
Hence $x_{-}\in\Sigma_{-}[v](y)$. A similar argument shows $x_{+}\in\Sigma_{-}[v](y)$. ∎
@@
If $u=T^{-}[v]$, $v\in SC(\mathbb{T})$ and the discrete standard map $F$ is locally Lipschitz continuous, then $\Sigma_{+}[u]$ is Lipschitz continuous.
@@
By Lemma 2.23, we know that $\Sigma_{+}[u]$ is Lipschitz continuous on $\Sigma_{-}[v](\mathbb{R})$. Let the Lipschitz constant of $\Sigma_{+}[u]$ on $\Sigma_{-}[v](\mathbb{R})$ be denoted by $K^{\prime\prime}$. For any $x_{1}\leq x_{2}\in\mathbb{R}$, by Lemma 3.4, there exist points $x_{1}^{\prime},x_{2}^{\prime}\in\Sigma_{-}[v](\mathbb{R})$ such that
$x_{1}\leq x_{1}^{\prime}\leq x_{2}^{\prime}\leq x_{2},\quad\Sigma_{+}[u](x_{1}^{\prime})=\Sigma_{+}[u](x_{1})\quad\text{ and }\quad\Sigma_{+}[u](x_{2}^{\prime})=\Sigma_{+}[u](x_{2}).$
Then,
$|\Sigma_{+}[u](x_{1})-\Sigma_{+}[u](x_{2})|=|\Sigma_{+}[u](x_{1}^{\prime})-\Sigma_{+}[u](x_{2}^{\prime})|.$
Since $x_{1}^{\prime},x_{2}^{\prime}\in\Sigma_{-}[v](\mathbb{R})$ and $\Sigma_{+}[u]$ is Lipschitz continuous on $\Sigma_{-}[v](\mathbb{R})$ with constant $K^{\prime\prime}$, we have
$|\Sigma_{+}[u](x_{1}^{\prime})-\Sigma_{+}[u](x_{2}^{\prime})|\leq K^{\prime\prime}|x_{1}^{\prime}-x_{2}^{\prime}|.$
From the relationship $x_{1}\leq x_{1}^{\prime}\leq x_{2}^{\prime}\leq x_{2}$, it follows that
$|x_{1}^{\prime}-x_{2}^{\prime}|\leq|x_{1}-x_{2}|.$
Therefore,
$|\Sigma_{+}[u](x_{1})-\Sigma_{+}[u](x_{2})|\leq K^{\prime\prime}|x_{1}-x_{2}|.$
This proves that $\Sigma_{+}[u]$ is Lipschitz continuous on the whole $\mathbb{R}$. ∎
3.3. $\Sigma_{+}[u]$ propagates singularities
We first give the following equivalent condition for singular points of functions in $T^{-}(SC(\mathbb{T}))$.
Proposition 3.6.
Let $u=T^{-}[v]$, where $v\in SC(\mathbb{T})$. Then $y\in\operatorname{Sing}(u)$ if and only if $(\Sigma_{+}[u])^{-1}(y)$ is a non-degenerate closed bounded interval.